3. A closed system initially containing 1.000 × 10-3 M H2 and 2.000 × 10-3 M I2 at 448℃ is allowed to reach equilibrium. At equilibrium, the HI concentration is 1.87 × 10-3 M. Given that the reaction equation is H2(g) + I2(g) ↔ 2 HI(g), find:
a. Equilibrium concentrations of the two reactants
b. Kc of the reaction taking place.
1
Expert's answer
2020-10-26T14:53:58-0400

"H_2(g) + I_2(g) \u2194 2 HI(g)"

initial concentration

"H_2 =1mM\\I_2=2mM"

"HI=0"


at equilibrium let's assume

"H_2 =(1-x)mM\\I_2=(2-x)mM\\HI=2x"


we are given equilibrium concentration of HI is 1.87mM

"2x=1.87\\x=0.935"


at equilibrium,

"H_2" ="1-0.935=0.065mM"

"I_2=2-0.935=1.065mM"

"HI=2\times0.935=1.87mM"


Kc of the reaction ="\frac{[HI]^2}{H_2I_2}=\frac{1.87\times1.87}{1.065\times0.065}=50.5"